OpenStudy (anonymous):

The ratio of the ages of Mandy and Sandy is 2:5. After 8 years, their ages will be in the ratio 1:2. What is the difference between their present ages?Caitlin says the difference in their ages is 24 years. In two or more complete sentences, explain whether or not Caitlin is correct. In your explanation, include the equations and calculations necessary to support or contradict Caitlin's answer.

OpenStudy (anonymous):

translate it to a set of equations, the first equation is 2m=5s and the second is m+8=2s+8 if you can solve that system of equations you should be able to find the answer

OpenStudy (anonymous):

would m/s= 2/5 (m+8)/ (5+8) = 1/2 work?